OpenStudy (anonymous):

Is x = 1 a solution of the equation 2 – 8x = –6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

plug in x = 1 and see if both sides are equal to each other 2-8(1) = -6 -6=-6

OpenStudy (anonymous):

2-8x=-6 -8x=-8 x=1

OpenStudy (anonymous):

^that works too